All of the following companies manufacture toasters. The following costs were incurred by each company during September:

Warner Co. Painted Trails Co. Catalina Inc. Sunrise Co.
Direct labor $11,900 $9,050 $72,100 $35,500
Rent 67,300 3,550 96,700 54,300
Utilities 2,600 1,300 16,200 7,800
Heating elements 86,800 41,900 100,800 64,700
Depreciation 5,800 9,600 27,900 37,700


Required:
1.
What is the total conversion cost for Warner Co.?
2. What is the total manufacturing overhead for Painted Trails Co.?
3. What is the total conversion cost for Catalina Inc.?
4. What is the total manufacturing overhead for Sunrise Co.?
Answer #1

1) Total Conversion Cost for Warner Co = 11900+67300+2600+86800+5800 = 174400

2) Total manufacturing overhead for Painted trails Co = 3550+1300+41900+9600 = 56350

3) Total Conversion Cost for Catalina Inc = 72100+96700+16200+100800+27900 = 313700

4) Total manufacturing overhead for Sunrise Co = 54300+7800+64700+37700 = 164500